# Stabilizer Entropy and entanglement complexity in the Sachdev-Ye-Kitaev model

### v2

@@ Abstract
Entanglement Spectrum Statistics (ESS).— Complex pattern of entanglement is characterized by universal properties of the statistics of the gaps in the entanglement spectrum, the so-called entanglement spectrum statistics (ESS) [19]. While chaotic systems, for instance, non-integrable systems obeying the eigenstate thermalization hypothesis (ETH) feature a universal, Wigner-Dyson (WD) behavior for the ESS, integrable, disordered free-fermion models feature a Poisson statistics in both the high energy states and the long time behavior away from equilibrium. Hybrid cases like MBL systems feature deviations from WD in polynomial time after a quantum quench [88]. The transition between the two regimes is due to the injection of non-stabilizer resources, which need to be scrambled around  [92, 81].
From the perspective of random matrix theory, if the eigenvalues of the reduced density matrix are uncorrelated, the distribution is Poissonian. Conversely, correlated eigenvalues follow the Wigner-Dyson universality class. This observation underscores that the entanglement patterns between the two interpolating regimes exhibit differing complexities. We focus on the probability density function (PDF) of the consecutive spacing ratios denoted as $P(r)$. To evaluate it we use the ascending eigenvalues $\{\lambda_{k}\}$ of the reduced density matrix and determine the spacing ratio as $r_{k}=\frac{\lambda_{k+1}-\lambda_{k}}{\lambda_{k}-\lambda_{k-1}}$, where $k=2,3,\dots,2^{R}-1$. The resulting ratios $\{r_{k}\}$ are plotted as a normalized histogram, excluding rare outliers with $r_{j}>10.0$ to ensure proper normalization and accurate binning [Odavic_Torre_Mijić_Davidović_Franchini_Giampaolo_2023, 61]. The explicit functional forms of the corresponding Poisson and Wigner-Dyson Gaussian ensembles are provided in  [5]. To complement this type of analysis, we also evaluate the averaged consecutive spacing ratio, defined as $\bar{r}=\langle\langle\frac{\min(s_{k,j},s_{k+1,j})}{\max(s_{k,j},s_{k+1,j})}\rangle\rangle_{2^{R}-2,M}$ where the spacings are given by $s_{k,j}=\lambda_{k+1,j}-\lambda_{k,j}$, and the index $j=1,2,\dots,M$ refers to the different reduced density matrices considered across ensemble or disorder realizations.
In Fig. 4, we compute the ESS $P(r)$ for different realizations of the GS of the $H_{g}$ model for several values of $g$. SYK4 ($g=0$) adheres to the universal WD distribution for GOE as it is expected for a real Hamiltonian. As soon as the model is perturbed by SYK2, the distribution starts getting closer to Poisson. In the lower right panel of Fig. 4, we show the behavior of the averaged consecutive spacing ratio $\bar{r}$ as a function of $g$, which shows a sudden jump as one moves from $g=0$. We can see that this feature of the entanglement complexity which is typical of chaotic systems is fragile in the model. Similar results hold for the MS, see [39].
@@
Figure 5: Ensemble averaged SRE $\mathcal{M}_{2}$ in the GS of the $H_{g}$. Panel (a): For all $g$ the ground states exhibit a non-vanishing amount of non-stabilizerness. Panel (b): The non-monotonicity can be easily captured by a grouping of points: green crosses for $N\!\!\mod 8=0$, blue for $N\!\!\mod 8=2,6$, and red circles for $N\!\!\mod 8=4$. The lines represent fits [39]. Panel (c): SRE without the normalization with the system size as compared to the Haar value. $|G\rangle\langle G|$ is a single qubit state which maximes SRE [39].
@@
Stabilizer Rényi Entropy (SRE)— Non-stabilizerness is an essential property for universal quantum computation [9]. To quantify this property, we use the Stabilizer Rényi Entropy (SRE) [51], defined as
$\mathcal{M}_{\alpha}(\Psi)=\dfrac{1}{1-\alpha}\log_{2}{\left(d^{-1}\sum_{P\in\mathcal{P}_{N}}|{\rm Tr}(\Psi P)|^{2\alpha}\right)}.$ (10)
@@
Here, $\Psi$ is the density matrix of a $N$-qubit state $|\Psi\rangle$, $d=2^{N}$ is the Hilbert space dimension, and $\mathcal{P}_{N}$ is the $N$-qubit Pauli group. The SREs for $\alpha\geq 2$ are good resource monotones [49]. Stabilizer entropies measure how far a state deviates from stabilizer states by analyzing its spread in the Pauli operator basis [51, 59]. The SRE stands out among non-stabilizerness monotones as it can be efficiently evaluated without the need for minimization procedures. In this work, we adopt the methodology outlined in Ref. [62] to compute the SRE. Specifically, we transform the ground state vectors obtained from exact diagonalization into matrix product state (MPS) tensor representations and employ the Perfect Sampling algorithm [36, 47, 23].
In Fig. 5, we show the numerical results for the scaling of SRE in the GS of the $H_{g}$ model. From panel $(a)$, we observe that the SYK-4 ($g=0$) case exhibits a higher degree of non-stabilizer resources compared to the SYK-2 ($g=1$) limit. From Fig. 5 ($a$), we observe that the SRE of the SYK-4 ground states does not exhibit a consistent monotonic increase with fermion number $N$, but instead shows a non-monotonic dependence. A more detailed analysis in Fig. 5 ($b$) reveals an oscillatory pattern in the SRE. The data is grouped and labeled according to the values of $N\!\!\mod 8$, as indicated in the captions. Each of the three groups corresponds to a distinct exponent of the damped exponential. This grouping is motivated by the observation in [24], where it was first noted that the SYK-4 Hamiltonian exhibits a particular particle-hole symmetry. The authors linked this symmetry to different Gaussian random matrix universality classes: GOE for $N\!\!\mod 8=0$, GUE for $N\!\!\mod 8=2,6$, and GSE for $N\!\!\mod 8=4$, which manifest in the Hamiltonian’s spectrum. We observe the effects of this symmetry specifically in the ground state, without considering the full Hamiltonian spectrum. In contrast to entanglement (see Fig. 2) and ground state energy [García-García_Verbaarschot_2016], the SRE appears to be much more sensitive to the finiteness of $N$. This highlights the unique ability of SRE to reveal hidden structures within many-body systems. A recent, striking example is the ability of SRE to detect a quantum phase transition that is not captured by entanglement in a system without a conventional order parameter [17].
@@
How does the average SRE evolve as the system size increases for the ground state of the SYK-4 Hamiltonian? This behavior is depicted in panel $(c)$ of Fig. 5. By fitting the data to a linear function [53], we obtain the following expression for the average SRE:
$\langle\mathcal{M}_{2}\rangle_{M}^{\rm fit}\sim-2.4+0.95\frac{N}{2}.$ (11)
@@
to leading order in the system size [51, 84, 62]. The factor $2$ in the denominator of the linear term arises because $N/2$ represents the support of the spin/qubit representation. Our results show that the SYK-4 ground states slightly deviate from the characteristics of fully quantum chaotic and universal states, with a linear prefactor difference of approximately $0.05$ when compared to Haar random states. A similar deviation was recently observed in the entanglement of middle-of-the-spectrum states in the SYK-4 model [38]. This provides evidence that low-temperature states, such as the ground states, fail to reach full universality.
Anti-Flatness.— A crucial point in the understanding of the relationship between SRE and entanglement comes from the realization that stabilizer states must have a flat entanglement spectrum [35]. Indeed, one can show that there is a strict relationship between the lack of flatness of the entanglement spectrum of the RDM and the SRE of the full state. In [80] it is shown that the linear SRE is exactly proportional to the average anti-flatness defined on a subsystem as $\mathcal{F}:={\rm Tr}[\rho^{3}_{\rm R}]-{\rm Tr}^{2}[\rho^{2}_{\rm R}]$. Another measure of anti-flatness is a numerically and analytically accessible quantity that is the logarithmic anti-flatness [62]
$\mathrm{F}(\rho_{R}):=2\left(S_{2}(\rho_{R})-S_{3}(\rho_{R})\right).$ (13)
In Supplementary Material [39], an excellent agreement across all system sizes is observed with the analytical results we derive for the SYK2 ($g=1$) limit where we obtained a closed-form expression for the logarithmic anti-flatness
$\mathrm{F}^{\rm SYK2}(R,f)=$ $2R(1-f)\sum\limits_{n=1}^{\infty}\dfrac{1}{n}\left(\dfrac{1}{2^{n}}-\dfrac{1}{2}\dfrac{3^{n}}{4^{n}}\right)$ (14)
×$\times\,_{2}F_{1}\left(\frac{1}{2},1-n\ ,2,4f(1-f)\right)$
where ${}_{2}F_{1}(a,b,c,d)$ is a hypergeometric function, $n>0$, and $20<f\leq 1/$ 2. For Haar random state in the large $N$ limit the logarithmic anti-flatness approach system-size-independent value of $\mathrm{F}^{\mathrm{Haar}}=\log{(5/4)}\approx 0.223$ [62].
@@
Its derivative with respect to the Rényi parameter at $\alpha=1$ is minus the variance of the entanglement Hamiltonian $H_{\rho}:=-\log\rho$ defined as
$-\operatorname{Var}_{\rho}(H_{\rho}):=-\langle\log^{2}\rho\rangle_{\rho}+\langle\log\rho\rangle_{\rho}^{2}=\partial_{\alpha}\tilde{S}_{\alpha}\Big{|}_{\alpha=1}$ (16)
It is also known as capacity of entanglement $C_{E}$ and it is yet another measure of anti-flatness [89, 75, 25]. This quantity is relevant because it quantifies the amount of non-local SRE, that is, the SRE that cannot be undone by local unitary operations [14]. For Haar random states, one can compute [25, 63]
$\partial_{\alpha}\tilde{S}_{\alpha}^{\rm Haar}\Big{|}_{\alpha=1}=\dfrac{11}{4}-\dfrac{\pi^{2}}{3}\approx-0.539868.$ (17)
@@
which we use in our analysis as a reference value.
@@
In Fig. 6, we compute the capacity of entanglement $C_{E}$ as a probe of non-local SRE as a function of $g$ in both the GS and MS of $H_{g}$. We see that SYK4 features a Haar-value capacity of entanglement, which is in perfect agreement in the MS. The finite-size scaling of the inset shows again the symmetric behavior of robustness observed earlier: while for the GS the SYK4 features are fragile, they are robust and can be extended all the way for any $g<1$ for the MS.
Figure 6: Ensemble averaged half-system capacity of entanglement in the GSs (left panel) and MSs (right panel) of $H_{g}$. The inset shows how the relative gap defined as $\Delta C_{\rm E}=|(C_{\rm E}-C_{\rm E}^{\rm Haar})/C_{\rm E}^{\rm Haar}|$ of $g=0$ (SYK4-) and $g=1$ (SYK2-) scales with respect to the Haar value. More on the $C_{\rm E}$ in the SYK-2 can be found in [7].
Conclusions and Outlook.— In this paper, we studied the interplay between entanglement and non-stabilizer resources in the SYK4+SYK2 model through several figures of merit that involve probes into the non-local character of stabilizer entropy, the entanglement spectrum statistics, and the adherence of such quantities to predictions from random matrix theory. We show that SYK4 features a complex pattern of this interplay while the integrable SYK2 is not universal. Moreover, we show the robustness of the universal features of SYK4 and find that they vary greatly between the ground state and highly excited states. For this reason, it would be interesting to study the behavior of such interplay away from equilibrium and its role in operator scrambling [79]. The quantum advantage of SYK batteries could be related to an optimal usage of both entanglement [3] and non stabilizer resources [15, 16, 31].
One of the most striking features of SYK4 is its dual JT gravity at low temperature. We have shown that SYK4 is fragile under a two-body pertrubation in the ground state while robust at high-energy, leaving open the question of its robustness at low temperature, which we plan to answer in a future work.
@@ Appendix A On the evaluation of Stabilizer Rényi Entropy
Here, we outline key details of the Perfect Sampling algorithm [9, 10], essential for obtaining reliable estimates of the Stabilizer Rényi Entropy (SRE) presented in Fig. 5. Starting from a state vector supported on $N/2$ qubits (where $N$ denotes the number of Majorana fermions) obtained via exact diagonalization, we represent it as a Matrix Product State (MPS) with a maximum bond dimension of $\chi=2^{N/2}$ and a truncation cutoff of $10^{-8}$ using the ITensor library [7, 6]. We then sample $10^{4}$ Pauli strings, leading to an absolute error in the stabilizer entropy estimation within the range $10^{-1}-10^{-2}$.
To verify the accuracy of this approach, we consider smaller system sizes in Fig. 7, where the SRE can be computed exactly from the state vector, evaluating all the $4^{N}$ expectation values of the Pauli strings, confirming that the method achieves sufficient precision. A similar approach was employed in [14]. The open-source ITensor implementation of this method will be made available in a forthcoming publication [18].
In Fig. 5 (b) the green crosses are related to sizes such that $N\!\!\mod 8=0$, for which the fit parameters are $a=0.57,b=0.036$, blue crosses for $N\!\!\mod 8=2,6$, with $a=0.44,b=0.09$, and red circles for $N\!\!\mod 8=4$, with $a=0.46,b=0.07$. The lines represent fits to the function $f(x)=a(1-\exp{(-bx)})$.
In Fig. 5(c), we highlight a factorizable $n$-qubit state with the maximal amount of single-qubit magic, shown in magenta. It is known that there is a single qubit state which maximizes the SRE, the golden state $|G\rangle\langle G|=\frac{1}{2}\left(I+\frac{X+Y+Z}{\sqrt{3}}\right)$. The stabilizer entropy of its $n$-qubits product state is $\mathcal{M}_{2}(|G\rangle\langle G|^{\otimes n})=n\log_{2}\left(\frac{3}{2}\right)$.

### v3

@@ III.3 Entanglement Spectrum Statistics (ESS)
Complex pattern of entanglement is characterized by universal properties of the statistics of the gaps in the entanglement spectrum, the so-called entanglement spectrum statistics (ESS) [48]. Chaotic systems, such as non-integrable quantum systems obeying the eigenstate thermalization hypothesis (ETH), feature a universal, Wigner-Dyson (WD) behavior for the ESS; while integrable, disordered free-fermion models feature a Poisson statistics in both the high energy states and the long time behavior away from equilibrium. Hybrid cases like MBL systems feature deviations from WD in polynomial time after a quantum quench [44]. The transition between the two regimes is due to the injection of non-stabilizer resources, which are scrambled around [42, 66]. From the perspective of random matrix theory, if the eigenvalues of the reduced density matrix are uncorrelated, the distribution is Poissonian. Conversely, correlated eigenvalues follow the Wigner-Dyson universality class.
Here we focus on the probability density function (PDF) of the consecutive spacing ratios denoted as $P(r)$. To evaluate it we use the ascending eigenvalues $\{\lambda_{k}\}$ of the RDM and determine the spacing ratio as
$r_{k}=\frac{\lambda_{k+1}-\lambda_{k}}{\lambda_{k}-\lambda_{k-1}},\quad k=2,3,\dots,2^{R}-1.$ (15)
The resulting ratios $\{r_{k}\}$ are plotted as a normalized histogram, excluding rare outliers with $r_{j}>10.0$ to ensure proper normalization and accurate binning [67, 68]. The explicit functional forms of the corresponding Poisson and Wigner-Dyson Gaussian ensembles are provided in  [69] and also Appendix E. To complement this type of analysis, we also evaluate the averaged consecutive spacing ratio, defined as
$\bar{r}=\Bigg{\langle}\Bigg{\langle}\frac{\min(s_{k,j},s_{k+1,j})}{\max(s_{k,j},s_{k+1,j})}\Bigg{\rangle}\Bigg{\rangle}_{2^{R}-2,M}.$ (16)
where the spacings are given by $s_{k,j}=\lambda_{k+1,j}-\lambda_{k,j}$, and the index $j=1,2,\dots,M$ refers to the different RDMs considered across ensemble or disorder realizations.
Figure 5: ESS of the GS RDM eigenvalues of the $H_{g}$ model for different values of $g$. We superimpose the analytical curves for the Wigner-Dyson (dashed) and Poisson (blue continuous) distribution for comparison. System size $N=22$ and number of realizations is $M=100$. The number of bins used for the histogram is 100. The colored numbers represent the KL divergence of data against known distributions described in Appendix E.
@@
Figure 6: Ensemble averaged SRE $\mathcal{M}_{2}$ in the GS of the $H(g)$. Panel (a): For all $g$ the ground states exhibit a non-vanishing amount of non-stabilizerness. Panel (b): The non-monotonicity can be easily captured by a grouping of points: green crosses for $N\!\!\mod 8=0$, blue for $N\!\!\mod 8=2,6$, and red circles for $N\!\!\mod 8=4$. The lines represent fits to the data. More details on the obtained fitting parameters are provided in Appendix A. Panel (c): SRE without the normalization with the system size as compared to the Haar value. $|G\rangle\langle G|$ is a single qubit state which maximizes SRE (see Appendix A), indicating beyond local non-stabilizerness present in the model as $N\to\infty$.
In Fig. 5, we compute the ESS $P(r)$ for different realizations of the GS of the $H(g)$ model for several values of $g$. The SYK-4 model adheres to the universal WD distribution for GOE. Once the model is perturbed by an SYK-2 term, the distribution shifts closer to a Poisson distribution. In the lower right panel of Fig. 5, we show the behavior of the averaged consecutive spacing ratio $\bar{r}$ as a function of $g$, which shows a sudden jump as one moves from $g=0$. We can see that this feature of the entanglement complexity, which is typical of chaotic systems, is fragile in the model. Similar results hold for the MS, see Appendix G.
III.4 Stabilizer Rényi Entropy
@@
Non-stabilizerness is an essential property for universal quantum computation [70]. To quantify this property, we used the so-called Stabilizer Rényi Entropy (SRE) $\mathcal{M}_{\alpha}$ [30],
$\mathcal{M}_{\alpha}(\Psi)=\dfrac{1}{1-\alpha}\log_{2}{\left(d^{-1}\sum_{P\in\mathcal{P}_{N}}|{\rm Tr}(\Psi P)|^{2\alpha}\right)},$ (17)
@@
the unique computable measure of non-stabilizerness [71]. The SRE can be efficiently evaluated without any minimization procedure. In Eq.17, $\Psi$ is the density operator of a $N$-qubit state $|\Psi\rangle$, $d=2^{N}$ is the Hilbert space dimension, and $\mathcal{P}_{N}$ is the $N$-qubit Pauli group. The SREs for $\alpha\geq 2$ are good resource monotones [71]. Stabilizer entropies measure how far a state deviates from stabilizer states by analyzing its spread in the Pauli operator basis [30, 72, 73].
[…]
@@
How does the average SRE evolve as the system size increases in the GS and MS of the SYK-4 Hamiltonian? This behavior is depicted in panel $(c)$ of Fig. 6 for the GS, and in the inset of Fig. 15 in Appendix H for the MS. By fitting the data to a linear function [79], we obtain the following expression for the average SRE:
$\mathcal{M}_{2}^{\rm SYK-4,GS}\sim-2.4+0.95\frac{N}{2}\;\ ,$ (18)
$\mathcal{M}_{2}^{\rm SYK-4,MS}\sim-2.6+0.96\frac{N}{2}\;\ .$ (19)
@@
to leading order in the system size [30, 80, 64]. The factor 222 in the denominator of the linear term arises because $N/2$ represents the support of the spin/qubit representation. Our results in Eq. (18) and Eq. (19), as compared to Eq. (20), indicate that both the GS and MS of the SYK-4 model exhibit deviations from the behavior expected of fully quantum chaotic and universal states. Specifically, we observe a linear prefactor difference of approximately 0.050.050.05 relative to Haar-random states. These observations support the conclusion that the eigenstates of the interpolated SYK-4 + SYK-2 model do not fully achieve universality.
III.5 Entanglement Spectrum Anti-Flatness
A crucial insight in understanding the relationship between non-stabilizerness and entanglement resources comes from the fact that stabilizer states have a flat entanglement spectrum [81]. Indeed, it is known that there is a strict relationship between the lack of flatness of the entanglement spectrum of a RDM and the non-stabilizerness of the full state. This is remarkable as anti-flatness is a local (that is, pertainind to the reduced density operator) quantity while SRE pertains to the full state. More specifically, it was shown [82] that there is a proportionality relation between the linear 2-SRE (see Eq. (17)), $\mathcal{M}_{2}^{\rm lin}(\Psi)=1-d^{-1}\sum_{P\in\mathcal{P}_{N}}|{\rm Tr}(\Psi P)|^{4}$, and a particular measure of the anti-flatness of a RDM. This establishes an explicit connection between bipartite entanglement properties and the total system’s non-stabilizerness response. In particular, a non-flat entanglement spectrum implies that the state is both entangled with respect to the chosen bipartition and possesses non-stabilizer features. Thus, anti-flatness serves as a diagnostic of magic that is invariant under local unitaries acting within each subsystem of the bipartition.
Figure 7: Absolute value of the averaged half-system capacity of entanglement $C_{E}(\rho_{\mathrm{R}})$ in the ground state (left panel) and middle-of-the-spectrum states (right panel) of $H_{g}$.
The non-local non-stabilizerness of a state is defined as the component of its non-stabilizerness that cannot be removed by local unitary operations [34, 38]. Thus, it is natural to define the non-local SRE to measure this quantity. The non-local SRE [34] of a state $|\Psi\rangle$ is defined as $\mathcal{M}_{\alpha}^{NL}(|\Psi\rangle)=\underset{U_{A}\otimes U_{B}}{\min}\mathcal{M}_{\alpha}(U_{A}\otimes U_{B}|\Psi\rangle)$, where the minimization is over all local unitaries acting on the bipartition $A\otimes B$. A vanishing value of $\mathcal{M}_{\alpha}^{NL}(|\Psi\rangle)$ means that the non-stabilizerness of the state is entirely local, thus it can be eliminated by local unitaries. In this case, the entanglement spectrum of the reduced density matrix is flat, establishing a direct connection between non-local non-stabilizerness and the structure of bipartite entanglement.
@@ III.5 Entanglement Spectrum Anti-Flatness
Its derivative, with respect to the Rényi parameter at $\alpha=1$, is minus the variance of the entanglement Hamiltonian $H_{\rho}:=-\log\rho$,
$-\operatorname{Var}_{\rho}(H_{\rho})=-\langle\log^{2}\rho\rangle_{\rho}+\langle\log\rho\rangle_{\rho}^{2}=\partial_{\alpha}\tilde{S}_{\alpha}\Big{|}_{\alpha=1}\ ;$ (22)
also known as capacity of entanglement $C_{E}$, and it is able to quantify the anti-flatness of a spectrum [84, 85, 86]. For Haar random states, one can compute
$C_{E}^{\rm Haar}=\partial_{\alpha}\tilde{S}_{\alpha}^{\rm Haar}\Big{|}_{\alpha=1}=\dfrac{11}{4}-\dfrac{\pi^{2}}{3}\approx-0.539868.$ (23)
@@
which we use in our analysis as a reference value [86, 87]. We note that $C_{\rm E}$ for the SYK-2 has been obtained in [88]
@@
In Fig. 7, we compute the averaged capacity of entanglement $C_{E}$ as a function of $g$ in both the GS and MS state of $H(g)$. We observe that SYK-4 features a Haar-like capacity of entanglement, which is in a good agreement in the MS. Our study of the capacity of entanglement supports our claims on robustness/fragility observed with previous probes: while for the GS, the SYK-4 features are fragile, they are robust and can be extended for any $g<1$ in the MS.
In Appendix D, we show the results for another measure of antiflatness, the logarithmic anti-flatness, expressed as the difference between Rényi entropies. For this measure we obtain an analytical expression of it incthe GS of the SYK-2 Hamiltonian. We find that the general behavior of this quantity strongly follows the trends of the capacity of entanglement shown here.
IV Conclusions and Outlook
In this work, we explored the interplay between entanglement and non-stabilizer resources in the interpolating SYK-4+SYK-2 model using a range of diagnostic tools. These include measures such as entanglement entropy, stabilizer entropy, entanglement spectrum statistics, and their comparison with predictions from random matrix theory and Haar-random states [80]. We find that, while several properties, especially in the middle of the spectrum, closely resemble Haar-random behavior, significant deviations persist. In particular, the stabilizer entropy systematically is different compared to the Haar expectation when extrapolated to the thermodynamic limit, both for ground state and for typical high-energy eigenstates. In contrast, the capacity of entanglement for high-energy states in SYK-4 aligns well with the Haar predictions, revealing a nuanced and measure-dependent picture of quantum complexity.
By introducing a Kullback-Leibler divergence-based probe of the entanglement spectrum, we identified a complexity transition in the GS of the interpolated SYK model. We find that the critical interpolation parameter scales as $\sim N^{-3/4}$ (where we set the coupling constant $J=1$), indicating the instability of the SYK-4 phase in the presence of arbitrarily small quadratic perturbations in the thermodynamic limit in the GS of the interpolated model. This result extends the findings of Ref. [14] and demonstrates that this information-theoretic probe is significantly more sensitive than the other diagnostics we employed.
Altogether, our study contributes to the ongoing discussion on the extent to which finite-sized SYK-4 model captures features of quantum black holes and fast scramblers [89]. An interesting future direction would be to study the behavior of stabilizer entropy and related diagnostics analytically in the double-scaled limit [52, 90], where greater theoretical control may offer deeper insights into the structure of quantum complexity in the SYK model.
The quantum advantage of SYK-based quantum batteries may stem from the optimal exploitation of both entanglement [91] and non-stabilizer resources [92, 93, 94], highlighting a promising avenue for future research as well.
One of the most remarkable aspects of SYK-4 is its duality with JT gravity in the low-temperature regime [7]. While our results show that SYK-4 is fragile under two-body perturbations in the GS but remains robust at high energies, the question of its stability at low temperatures remains open—an issue we aim to address in future work.
@@ Appendix A On the evaluation of Stabilizer Rényi Entropy
Here, we outline key details of the Perfect Sampling algorithm [74, 75], essential for obtaining reliable estimates of the Stabilizer Rényi Entropy (SRE) presented in Fig. (5). Starting from a state vector supported on $N/2$ qubits (where $N$ denotes the number of Majorana fermions) obtained via exact diagonalization, we represent it as a Matrix Product State (MPS) with a maximum bond dimension of $\chi=2^{N/2}$ and a truncation cutoff of $10^{-8}$ using the ITensor library [95, 96]. We then sample $10^{4}$ Pauli strings, leading to an absolute error in the stabilizer entropy estimation within the range $10^{-1}-10^{-2}$.
To verify the accuracy of this approach, we consider smaller system sizes in Fig. 8, where the SRE can be computed exactly from the state vector, evaluating all the $4^{N}$ expectation values of the Pauli strings, confirming that the method achieves sufficient precision. A similar approach was employed in [64].
In Fig. (5b) the green crosses are related to sizes such that $N\!\!\mod 8=0$, for which the fit parameters are $a=0.57,b=0.036$, blue crosses for $N\!\!\mod 8=2,6$, with $a=0.44,b=0.09$, and red circles for $N\!\!\mod 8=4$, with $a=0.46,b=0.07$. The lines represent fits to the function $f(x)=a(1-\exp{(-bx)})$.
In Fig. (5c), we highlight a factorizable $n$-qubit state with the maximal amount of single-qubit magic, shown in magenta. It is known that there is a single qubit state that maximizes the SRE, the golden state $|G\rangle\langle G|=\frac{1}{2}\left(I+\frac{X+Y+Z}{\sqrt{3}}\right)$. The stabilizer entropy of its $n$-qubits product state is $\mathcal{M}_{2}(|G\rangle\langle G|^{\otimes n})=n\log_{2}\left(\frac{3}{2}\right)$.